OpenStudy (anonymous):

Solve the equation 3tan(2x+15)=4 for 0

OpenStudy (anonymous):

3tan(2x+15)=4 tan(2x+15)=4/3 2x+15=arctan(4/3) 2x =arctan(4/3)-15 x =(arctan(4/3)-15)/2

OpenStudy (anonymous):

Pick the value that is between 0 and 180

OpenStudy (anonymous):

Is that the only value?

OpenStudy (anonymous):

There are multiple values because the equation is ______ I forgot the word but when you graph a cos and sin you just see it go on and on and on. So you can have multiple values for x but they want you to find it for a very specific spot between 0 and 180

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3tan%282x%2B15%29%3D4+solve+for+x See the graph

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Usually they ask you in terms of pi it should be between 0 and pi And the word is periodic and the period is between 0 and 2pi